40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sun Feb 28 16:59:39 2021, rana, gautam, Update, LSC, more PRMI checks here: what it is ain't exactly clear RAMestimate.png
    Reply  Sun Feb 28 22:53:22 2021, gautam, Update, LSC, more PRMI checks here: what it is ain't exactly clear PRMI_SBres_REFL55.pngPRMI1f_noArmssensMat.pdfPRMI1f_noArmssensMat.pdfPRMI_locked.pngactTFs.pdf
Message ID: 15850     Entry time: Sun Feb 28 22:53:22 2021     In reply to: 15849
Author: gautam 
Type: Update 
Category: LSC 
Subject: more PRMI checks here: what it is ain't exactly clear 

I looked into this a bit more and crossed off some of the points Rana listed. In order to use REFL 55 as a sensor, I had to fix the frequent saturations seen in the MICH signals, at the nominal (flat) whitening gain of +18 dB. The light level on the REFL55 photodiode (13 mW), its transimpedance (400 ohm), and this +18dB (~ x8) gain, cannot explain signal saturation (0.7A/W * 400 V/A * 8 ~ 2.2kV/W, and the PRCL PDH fringe should be ~1 MW/m, so the PDH fringe across the 4nm linewidth of the PRC should only be a couple of volts). Could be some weird effect of the quad LT1125. Anyway, the fix that has worked in the past, and also this time, is detailed here. Note that the anomalously high noise of the REFL55_Q channel in particular remains a problem. After taking care of that, I did the following:

  1. PRMI (ETMs misaligned) locking with sidebands resonant in the PRC was restored - REFL55_I was used for PRCL sensing and REFL55_Q was used for MICH sensing. The locks are acquired nearly instantaneously if the alignment is good, and they are pretty robust, see Attachment #1 (the lock losses were IMC related and not really any PRC/MICH problem).
  2. Measured the loop OLTFs using the usual IN1/IN2 technique. The PRCL loop looks just fine, but the MICH loop UGF is very low apparently. I can't just raise the loop gain because of the feature at ~600 Hz. Not sure what the origin of this is, it isn't present in the analogous TF measurement when the PRMI is locked with carrier resonant (REFL11_I for PRCL sensing, AS55_Q for MICH sensing). I will post the loop breakdown later. 
  3. Re-confirmed that the MICH-->PRCL coupling couldn't be nulled completely in this config either.
    • The effect is a geometric one - then 1 unit change in MICH causes a 1/sqrt(2) change in PRCL. 
    • The actual matrix element that best nulls a MICH drive in the PRCL error point is -0.34 (this has not changed from the PRMI resonant on carrier locking). Why should it be that we can't null this element, if the mechanical transfer functions (see next point) are okay?
  4. Looked at the mechanical actuator TFs are again (since we forgot to save plots on Friday), by driving the BS and PRM with sine waves (311.1 Hz), one at a time, and looking at the response in REFL55_I and REFL55_Q. Some evidence of some funkiness here already. I can't find any configuration of digital demod phase that gives me a PRCL/MICH sensing ratio of ~100 in REFL55_I, and simultaneously, a MICH/PRCL sensing ratio of ~100 in REFL55_Q. The results are in Attachments #5
  5. Drove single frequency lines in MICH and PRCL at 311.1 and 313.35 Hz respectively, for 5 minutes, and made the radar plots in Attachments #2 and #3. Long story short - even in the "nominal" configuration where the sidebands are resonant in the PRC and the carrier is rejected, there is poor separation in sensing. 
    • Attachments #2 is with the digital REFL55 demod phase set to 35 degrees - I thought this gave the best PRCL sensing in REFL55_I (eyeballed it roughly by looking at ndscope free-swinging PDH fringes).
    • But the test detailed in bullet #4, and Attachments #2 itself, suggested that PRCL was actually being sensed almost entirely in the Q phase signal.
    • So I changed the digital demod phase to -30 degrees (did a more quantitative estimate with free-swinging PDH fringes on ndscope, horn-to-horn voltages etc).
    • The same procedure of sine-wave-driving now yields Attachments #3. Indeed, now PRCL is sensed almost perfectly in REFL55_I, but the MICH signal is also nearly in REFL55_I. How can the lock be so robust if this is really true? 
  6. Attachments #4 shows some relevant time domain signals in the PRMI lock with the sidebands resonant. 
    • REFL11_I hovers around 0 when REFL55_I is used to sense and lock PRCL - good. The m/ct calibration for REFL11_I and REFL55_I are different so this plot doesn't directly tell us how good the PRCL loop is based on the out-of-loop REFL11_I sensor.
    • ASDC is nearly 0, good.
    • POP22_I is ~200cts (and POP22_Q is nearly 0) - I didn't see any peak at the drive frequency when driving PRCL with a sine wave, so no linear coupling of PRCL to the f1 sideband buildup, which would suggest there is no PRCL offset.
    • Couldn't do the analogous test for AS110 as I removed that photodiode for the AS WFS - it is pretty simple to re-install it, but the ASDC level already doesn't suggest anything crazy here.

Rana also suggested checking if the digital demod phase that senses MICH in REFL55_Q changes from free-swinging Michelson (PRM misaligned), to PRMI aligned - we can quantify any macroscopic length mismatch in the PRC length using this measurement. I couldn't see any MICH signal in REFL55_Q with the PRM misaligned and the Michelson fringing. Could be that +18dB is insufficient whitening gain, but I ran out of time this afternoon, so I'll check later. But not sure if the double attenuation by the PRM makes this impossible.

Attachment 1: PRMI_SBres_REFL55.png  30 kB  | Hide | Hide all
PRMI_SBres_REFL55.png
Attachment 2: PRMI1f_noArmssensMat.pdf  147 kB  | Hide | Hide all
PRMI1f_noArmssensMat.pdf
Attachment 3: PRMI1f_noArmssensMat.pdf  145 kB  | Hide | Hide all
PRMI1f_noArmssensMat.pdf
Attachment 4: PRMI_locked.png  155 kB  Uploaded Mon Mar 1 11:01:00 2021  | Hide | Hide all
PRMI_locked.png
Attachment 5: actTFs.pdf  246 kB  Uploaded Mon Mar 1 11:40:10 2021  | Hide | Hide all
actTFs.pdf
ELOG V3.1.3-